Sam paid $12,000 of his salary as taxes this year. What he paid this year as taxes is 20% more than what he paid last year. If he paid 1/4 of his salary as taxes last year, what was his salary last year?
Thanks for taking time to answer my quastion.

Answer by jorel555(1290)

12000 equals taxes this year. 20% more than last year is:
1.2n=12000
n=12000/1.2=10000
1/4 salary=10000
1 salary=$40000 earned by Sam last year!!!!!!!!!!!!!!!!!!